OpenStudy (anonymous):

3. Find the measure of an angle between 0º and 360º coterminal with an angle of –110º in standard position. 250º 20º 110º 70º

OpenStudy (mathstudent55):

If you add or subtract multiples of 360 to an angle in standard position, you get coterminal angles.

OpenStudy (anonymous):

2,180 3,120 4,90 5,72

OpenStudy (mathstudent55):

?

OpenStudy (anonymous):

those are multiplys that equal 360

OpenStudy (mathstudent55):

Multiples of 360: 360 * 0 = 0 360 * 1 = 360 360 * 2 = 720 etc. Now add each multiple to -110 until you find the multiple that gives a sum between 0 and 360. -110 + 0 = -110 not between 0 and 360 -110 + 360 =?

OpenStudy (anonymous):

250

OpenStudy (mathstudent55):

correct